This one is giving me alot of trouble. The question is: Given f(x)=3x+5, find f(a-1). Please help explain this to me!!
Notice that x is an independent variable which can be replaced with any number; numbers; or other variables...in this case we want to substitute it with (a-1)...
f(x)=3x+5 then
f(a-1)=3(a-1)+5....any time you substitute x with anything you must place that in parenthesis like I did to show the substitution
so
f(a-1)=3a-3+5 (by distributing 3 to (a-1), i.e. multiplying 3 with variables inside the parenthesis to get rid of parenthesis), then
f(a-1)=3a+2 by simplifying (adding like terms together)

Found 2 solutions by eperette, Alan3354:
Answer by eperette(173)
Answer by Alan3354(21583)

This one is giving me alot of trouble. The question is: Given f(x)=3x+5, find f(a-1). Please help explain this to me!!
-----------------
Notice that x is an independent variable which can be replaced with any number; numbers; or other variables...in this case we want to substitute it with (a-1)...
f(x)=3x+5 then
f(a-1)=3(a-1)+5....any time you substitute x with anything you must place that in parenthesis like I did to show the substitution
so
f(a-1)=3a-3+5 (by distributing 3 to (a-1), i.e. multiplying 3 with variables inside the parenthesis to get rid of parenthesis), then
f(a-1)=3a+2 by simplifying (adding like terms together)
-------------------
That's all there is to it.